# Post-detection inference for sequential changepoint localization

### v3

@@ 4.3 Adaptive confidence set for parametric pre- and post-change classes
[…]
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$. Note that the sequences must be independent as $j$ varies. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section C of the Supplementary. We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$. Similar to (4.4), the adaptive confidence set is defined as
@@
where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Algorithm 3 contains an overview. If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.3 and provides coverage guarantees, as elaborated in Section F of the Supplementary.
As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section E of the Supplementary. Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative. We provide the details in Section C of the Supplementary, along with two concrete examples.
@@
Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.5), for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$. For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.3 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
@@
As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$. So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.4 and 4.5 reduces to Theorem 4.1.
@@ 5 Confidence set for pre- and post-change parameters
Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data ${X_{1},\ldots,X_{t-1}}$, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$. The final confidence set for $\theta_{0}$ is the following:
$\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ (5.2)
We formalize its coverage guarantee in the next theorem, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods [39, 36].
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{1}.$
@@ 6 Experiments
We validate our methods across various settings and demonstrate their practicality.
Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter. We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$. After detection, we construct both the proposed confidence sets (universal (3.6) and adaptive (4.1)), with $B=N=100$, $\alpha=0.05$. Fig. 6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across 555 runs. Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across 500500500 runs.
Figure 6.1: Setting I: The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$. The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$. Results of 555 independent simulations are shown.
@@
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$. After detection, we construct both the proposed confidence sets for $T$ (universal (3.6) with $\alpha=0.075$ and adaptive (4.4) [^6] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section B.1 (and Section C.1) of the Supplementary. with $\alpha=0.05,\beta=0.025$) with target conditional coverage 0.9250.9250.925. For implementing (3.6) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where ${\theta_{1}^{*}}$ is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above. In Table 6.4, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across 500500500 runs. Fig. 6.3 visualizes the confidence sets (3.6) and point estimates (B.2) across 555 random runs.
Figure 6.3: Setting II: First $T-1$ samples are drawn from $N(0,1)$ and the rest from $N(1,1)$. The point estimates (B.2) are shown in a vertical red dashed line, and confidence sets (universal (3.6)) are shown in red points. $F_{0}=N(0,1)$, $\mathcal{P}_{1}=\{N(\mu,1):\mu\geq 0.9\}$, $N=100$, $\alpha=0.075$. Results of 555 independent simulations are shown.
@@
100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.21 25.72 3.95 16.87
100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 17.85 24.11 3.67 16.21
500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.89 30.83 3.45 17.63
500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.14 28.56 3.48 15.81
@@
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$. After detection, we construct both the proposed confidence sets (universal (3.8) with $\alpha=0.1$ and adaptive (4.5) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage 0.90.90.9. For implementing (3.8), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II. $\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space. Fig. 6.5 visualizes the confidence sets (3.8) and point estimates (C.2) across 5 random runs. Table 6.6 reports the results averaged across 500500500 runs.
Figure 6.5: Setting III: First $T-1$ samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1)$. The point estimates (C.2) are shown in a vertical red dashed line. Confidence sets (universal (3.8)) are shown in red points. $\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=100$, $\alpha=0.1$. Results of 555 random simulations are shown.
@@
100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.91 26.84 4.36 25.81
100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 18.63 25.71 4.19 23.15
500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.12 31.45 4.03 24.48
500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.86 30.66 4.07 23.07
Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on [0,1][0,1][ 0 , 1 ]) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$. For detection, we use an e-detector [26] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$. Post-detection, we construct the confidence sets (3.6). The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [15] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [19]. We divide the interval [0,1][0,1][ 0 , 1 ] into 101010 equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation. We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$. We use the point estimator
@@
Table 6.7 contains average results across 500500500 runs and Fig. 6.8 visualizes across 555 runs.
[…]
@@
[…]
Figure 6.8: Setting IV: First $T-1$ samples are drawn from $U(0,1)$ and the remaining observations from the density $f(x)=4(1-x)^{3},x\in[0,1]$. The point estimate (6.5) is shown in a vertical red dashed line, and the confidence set (3.8) is shown in red points. Results of 555 random simulations are shown.
[…]
@@
with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.3 with $F_{0}^{*}=\text{N}(0.5,1)$. After detection, we construct the confidence sets (3.8). The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{0.5,\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{\mu_{i}X_{i}-\mu_{i}^{2}/2}$, where $\mu_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}-0.5$ (see, e.g., Ch. 6.7.5 of [21]). We use the following point estimator:
@@
Figure 6.9: Setting V. First $T-1$ samples are drawn from $N(1,1)$ and the remaining samples are from $U(-1.2,0.8)$. Vertical red dashed lines show the point estimates (6.6). The confidence set (3.8) is marked in red points. Results of 555 independent simulations are shown.

### v4

@@ 4.3 Adaptive confidence set for parametric pre- and post-change classes
Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively. Instead of point estimates, we rely on interval estimates to achieve a predetermined control over the error. Suppose that we have some confidence interval (CI) procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$ that constructs CI using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{0}$ with coverage 1c1-c and a confidence sequence (CS) procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, that constructs CS using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{1}$ having coverage 1c1-c. We construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$, where $r_{t}^{*}$ is computed under Assumption 3.5.
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$. Note that the sequences must be independent as $j$ varies. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section D of the Supplementary. We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$. Similar to (4.3), the adaptive confidence set is defined as
@@
where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Algorithm 3 contains an overview. If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.5 and provides coverage guarantees, as elaborated in Section G of the Supplementary.
As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section F of the Supplementary. Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative. We provide the details in Section D of the Supplementary, along with two concrete examples.
@@
Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.4), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{\theta_{0}^{*},\infty}[\tau\geq t]$, for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.5 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
@@
As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$. So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.3 and 4.4 reduces to Theorem 4.1.
@@ 5 Confidence set for pre- and post-change parameters
Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data $X_{1},\ldots,X_{t-1}$}, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$. The final confidence set for $\theta_{0}$ is $\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ Next, we formalize its coverage guarantee, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods.
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{1}.$
@@ 6 Experiments
Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter. We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$. After detection, we construct both the proposed confidence sets (universal (3.5) and adaptive (4.1)), with $B=N=100$, $\alpha=0.05$. Fig. 6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across $5$ runs. Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across $500$ runs.
Figure 6.1: Setting I: The first T1T-1 observations are drawn from $N(0,1)$ and the rest from $N(1,1)$. The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
@@
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$. After detection, we construct both the proposed confidence sets for $T$ (universal (3.5) with $\alpha=0.075$ and adaptive (4.3) [^5] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section C.1 (and Section D.1) of the Supplementary. with $\alpha=0.05,\beta=0.025$) with target conditional coverage $0.925$. For implementing (3.5) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where $\theta_{1}^{*}$} is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above. In Table 6.3, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across $500$ runs. Fig. 6.4(a) shows the confidence sets (3.5) and point estimates (S2) across $5$ random runs.
@@
100 $0.75,\infty)$ 0.98 0.97 22.21 25.72 3.95 16.87
100 $0.9,\infty)$ 0.97 0.97 17.85 24.11 3.67 16.21
500 $0.75,\infty)$ 0.98 0.97 22.89 30.83 3.45 17.63
500 $0.9,\infty)$ 0.97 0.98 18.14 28.56 3.48 15.81
@@
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$. After detection, we construct both the proposed confidence sets (universal (3.9) with $\alpha=0.1$ and adaptive (4.4) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage $0.9$. For implementing (3.9), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II. $\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space. Fig. 6.4(b) visualizes the confidence sets (3.9) and point estimates (S2) across 5 random runs. Table 6.5 reports the results averaged across $500$ runs.
Figure 6.4: First T1T-1 samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1),T=100$. The point estimates (S2) are shown in a vertical red dashed line. Confidence sets (universal (3.9)) are shown in red points. $N=100$, $\alpha=0.1$. Results of $5$ random simulations are shown.
@@
100 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.91 26.84 4.36 25.81
100 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.97 18.63 25.71 4.19 23.15
500 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.12 31.45 4.03 24.48
500 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.98 18.86 30.66 4.07 23.07
Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on $[0,1]$) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$. For detection, we use an e-detector [28] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$. Post-detection, we construct the confidence sets (3.5). The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [17] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [20]. We divide the interval $[0,1]$ into $10$ equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation. We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$. We use the point estimator
@@
Table 6.6 shows average results across $500$ runs and Fig. 6.7(a) visualizes across $5$ runs.
Table 6.6: Setting IV: known pre-change ($\text{U}[0,1]$) but nonparametric post-change.
@@
500 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.99 0.93 21.27 2.94 15.78
Composite parametric pre- and nonparametric post-change (Setting V): Here, we consider the true (unknown) pre- and post-change distributions to be $\text{N}(1,1)$ and $\text{U}[-1.2,0.8]$ respectively. Let the pre-change class be $\mathcal{P}_{0}=\left\{\text{N}(\theta,1):\theta\geq 0.5\right\},$ and the post-change class be the set of all SubGaussians with a nonpositive mean: $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[e^{\lambda X}]\leq e^{-\lambda^{2}/2},\forall\lambda\in[0,\infty)\right\},$ $\mathcal{M}$ is the set of all probability measures on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. We use an e-detector [28],
@@
with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.5 with $P_{0}^{*}=\text{N}(0.5,1)$. After detection, we construct the confidence sets (3.9). The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{0.5,\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{\mu_{i}X_{i}-\mu_{i}^{2}/2}$, where $\mu_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}-0.5$ (see, e.g., Ch. 6.7.5 of [22]). We use the following point estimator:
@@
Figure 6.7: Vertical red dashed lines show the point estimates (6.6). The confidence set (3.9) is marked in red points. Results of $5$ independent simulations are shown.